OpenStudy (kearsten):

fan and medal Ralph has a piece of shelving 3 m long. He is going to cut it into 4 equal-length pieces. How long will each piece be?

OpenStudy (wolfman77):

I can help, if you like...

OpenStudy (kearsten):

thnx

OpenStudy (wolfman77):

Mhmm. No Problem.

OpenStudy (wolfman77):

Okay so, I am SO not good at explaining, but the answer is: Each piece will be about 2 feet 3 inches long. Do you need an explanation?......

OpenStudy (kearsten):

in fraction?

OpenStudy (wolfman77):

I don't think so.... Does it need to be a fraction?

OpenStudy (kearsten):

yes... if possible...

OpenStudy (wolfman77):

Okay, one sec.

OpenStudy (wolfman77):

I actually have no idea how to convert the foot/inches into a fraction. It's been awhile since I've done this. Is there an answer choices?

OpenStudy (kearsten):

no

OpenStudy (kevin):

3/4 :)

OpenStudy (sshayer):

3m=3*100=300 cm each piece=300/4=75 cm

OpenStudy (mathmale):

If the length of the whole plank is given in metric units (3 m), then the same unit applies to the length of each of the 4 pieces into with the 3 m plank is sawed. So you do NOT have to do any conversion. Rather, you need to divide 3 m by 4. Show your work and include the correct unit of measurement.